OpenStudy (anonymous):

Find the derivative of f(x) = 8/x at x = -1.

OpenStudy (anonymous):

@mebs

OpenStudy (anonymous):

i did what you showed me before and got 8

OpenStudy (psymon):

Almost.

OpenStudy (anonymous):

-8?

OpenStudy (psymon):

^_^

OpenStudy (anonymous):

Do what I did before. bring that x up and just the first principle and than simplify.\[step 1) f(x) = 8(x)^{-1}\] \[Step 2) f'(x) = 8(-1)x^{-1 -1}\] \[Step 3) f'(x) = \frac{-8}{x^{2}}\]

OpenStudy (anonymous):

awesome, i got it, -8

OpenStudy (anonymous):

Good. Just that step process until you can do it in the blink of an eye.

OpenStudy (anonymous):

ok awesome, thanks